OpenStudy (anonymous):

If a line with a slope of 4 crosses the y-axis at (0, 1), what is the equation of the line? y − 4x = 1 y + 4x = 1 x − 4y = 1 x + 4y = 1 Done

OpenStudy (jdoe0001):

http://www.algebra-class.com/images/form.gif

OpenStudy (texaschic101):

you will have to use the point slope formula : y - y1 = m(x - x1) set 1 (0,1).....where x1 = 0 and y1 = 1 slope (m) = 4 now sub your info into the formula... y - y1 = m(x - x1) y - 1 = 4(x - 0) (distribute through the parenthesis) y - 1 = 4x - 0 y - 4x = 1 <-- answer....its the first one